OpenStudy (anonymous):

Solve: Cos2θ = -√3/2

OpenStudy (anonymous):

cmon brainshot plz help me :)

OpenStudy (anonymous):

huh? @@!

OpenStudy (anonymous):

in the book it have 3 formulas

OpenStudy (anonymous):

cos 2A = cos^2A - sin^2A

OpenStudy (anonymous):

2cos^2A - 1 and 1 - 2 sin^2A

OpenStudy (anonymous):

Ha wow, I'm trying to do two things at once, sorry. I'm thinking of this as a calc prob. But back on topic, Cos(30 degrees) = √3/2 and we know that roots can be both negative and positive, right? This means that -√3/2 is a possibility, so it is 30/2 which is 15 degrees.

OpenStudy (anonymous):

where 30 degrees coming from ?

OpenStudy (anonymous):

Unit circle

OpenStudy (anonymous):

and it is 30 degrees because you look at the √3/2 ?

OpenStudy (anonymous):

15 degrees is not the answer though :|

OpenStudy (anonymous):

The answer are : 75 degrees, 105 degrees, 255 degrees and 285 degrees

OpenStudy (anonymous):

but how you get that :|

OpenStudy (anonymous):

Hi, yes, ahh I don't really remember this stuff. Isn't there as specific formula for this type of equation? I said 15 since I know that Cos30 = Root3 / 2, so I thought it would be 30 /2 as you are multiplying by two here, but that is incorrect because this is a circle out of 360 degrees, not 100, so you can't just divide by 2.

OpenStudy (anonymous):

not sure what formulas your trying to say here :|

OpenStudy (anonymous):

those are the 3 formulas that the book gave me

OpenStudy (anonymous):

this is gonna be one hell of a test tomorrow :|

OpenStudy (anonymous):

Now I feel guilty. Don't worry, I'll get someone to help you and while I do that, i'll try to pull up an old formula sheet with the formula for this on it.

OpenStudy (anonymous):

Oh no it's fine, I'm sorry for saying that lol :) It just that, this stuff is really confusing, I'm glad I got someone to help me you know :) so thanks for that haha :D

OpenStudy (anonymous):

I actually did manage to find the formula sheet, but it seems as though I was thinking of the Cos(a+b) formulas which do not help. But, I typed it into my calculator and I actually did get 15 as being equal to theta here. Are you sure it isn't correct??

OpenStudy (anonymous):

Wait is the equation Cos2x =-Root(3) / 2 or -Root (3/2) I've been going off the basis that it has been -Root(3) / 2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cos2x =-Root(3) / 2

OpenStudy (anonymous):

I'm sorry about that !!

OpenStudy (anonymous):

-Root (3/2) is the correct one sorry !

OpenStudy (anonymous):